# A Survey on Memory-Efficient Large-Scale Model Training in AI for Science

### v1

@@ III-B Mixed Precision Training
Mixed Precision Training is a deep learning technique that leverages data types of varying precisions to enhance training speed, reduce memory consumption, and decrease computational costs, all while preserving comparable accuracy. In mixed precision training introduced by Micikevicius et al. [78], model parameters, activations, and gradients are stored using FP16, while maintaining a copy of the FP32 parameters. The FP16 data type utilizes half the memory space of FP32 and offers improved computation speed. Nevertheless, FP16 has a reduced numerical range and accuracy compared to FP32, which may lead to numerical stability issues during training. To cope with it, FP16 parameters are used for forward and backward calculations, while FP32 parameters are used for model updates. Additionally, to mitigate the risk of gradient underflow during backpropagation, loss scaling is commonly applied, which helps to minimize model performance degradation. The mixed precision training process is illustrated in Fig. 5.
@@
While it needs to save an additional FP32 model parameter, the memory savings achieved by utilizing FP16 activations in mixed precision training surpass the memory costs associated with these extra FP32 parameters, leading to an overall memory usage reduction.
@@
Modern GPUs, such as NVIDIA’s Volta, Turing, and Ampere architectures, feature specialized Tensor Cores designed to accelerate mixed-precision matrix operations. However, BF16 support is limited to certain hardware platforms, such as Ampere and Hopper architectures. Since the Hopper architecture, FP8 Tensor Cores were added allowing for FP8 mixed precision training, which can further reduce memory consumption and accelerate computation [104], [105].
@@
Mixed precision training is a crucial technique for minimizing memory usage in training LLMs; however, it requires careful management to prevent numerical issues. As hardware and software continue to evolve, the methods and tools for mixed precision training are also advancing, facilitating the training of larger and more intricate models.
III-C Gradient Checkpointing
@@
In the training process of DNNs, the output of each layer (activations) needs to be stored in memory to calculate gradients during backpropagation. When dealing with large-scale models, long input sequences, or large batch sizes, the memory overhead caused by activations can become substantial.
The core idea behind gradient checkpointing is to trade computation for memory. This approach reduces memory usage by retaining only the intermediate activations of certain layers during forward propagation while recalculating the discarded activations when needed during backpropagation. The preserved intermediate activations are referred to as “checkpoints” [79].
@@
The number of parameters and computational complexity of various components in DNNs often vary. Korthikanti et al. [90] suggest that for transformer models, it is unnecessary to set checkpoints for the entire transformer layer for recomputation. Instead, selective recomputation can be performed by identifying components of the transformer layer where activations cost substantial memory yet require less computation. By setting checkpoints and recalculating selectively, it is possible to reduce memory overhead while minimizing computation.
III-D Offloading
SwapAdvisor [106] takes the data flow graph of a model as input and selects legitimate memory allocation and operator scheduling based on the graph. By precisely planning which and when tensors require swapping in and out, SwapAdvisor achieves optimal overlap between computation and communication. The integration of genetic algorithms enables the exploration of the vast search space for finding the optimal combination of memory allocation and operator scheduling. In contrast to previous heuristic-based swapping methods [107], [108], SwapAdvisor significantly enhances memory efficiency and computational performance through its dual optimization approach. Notably, due to the lack of consideration of GPU communication, SwapAdvisor has limitations in multi-GPU training scenarios.
ZeRO-infinity [110] enhances the capabilities of ZeRO-3 by adding support for heterogeneous memory, thereby leveraging the entire spectrum of memory types within the system (GPU, CPU, and SSD) to enable large-scale model training. By employing the Infinity Offload Engine to offload model states to NVMe storage and efficient communication optimizations, ZeRO-infinity enables superlinear scaling of training throughput. On 32 Nvidia V100 DGX-2 nodes, ZeRO-infinity supports Transformer models with up to 32T parameters, which is 50× the maximum size accommodated by 3D Parallelism, which is around 650B parameters.
III-E Gradient Accumulation
Gradient accumulation is an effective technique for training scenarios with limited memory and computing resources. It allows training models with a larger effective batch size when GPU memory is limited. The fundamental idea behind gradient accumulation is to gradually accumulate gradients from multiple small batches until a desired batch size is achieved, at which point a weight update is performed. For instance, if the total batch size we aim to use is 32, which exceeds the GPU memory capacity, we can divide it into smaller batches with batch sizes of 8, and perform 4 training iterations before conducting a weight update.
Algorithm 1 presents the pseudocode of gradient accumulation, outlining the implementation of this technique in the training process. Gradient accumulation is frequently utilized in hardware environments with limited resources and serves as an important optimization method in deep learning training.
@@ III-E Gradient Accumulation
III-F Memory Efficient Optimizers
Currently, adaptive optimizers like Adam [111] and AdamW [112] have nearly become the default optimizers for training LLMs. While Adam demonstrates impressive performance, it incurs a high memory cost due to the store of its optimizer states, including first-moment and second-moment estimates of gradients. This results in a memory overhead that is at least double the size of the model itself, which poses a significant challenge during pre-training. To support such memory-intensive algorithms, techniques like offloading and sharding are often enabled in practice; however, these methods can lead to increased training latency and decreased efficiency. Consequently, there is a strong interest in developing optimizers that utilize less memory while maintaining efficiency.
Shazeer et al. introduced a memory-efficient adaptive optimizer known as Adafactor [113], which significantly lowers memory requirements while delivering training performance comparable to Adam. Assuming the model parameter matrix is denoted as $W\in R^{m\times n}$, by applying low-rank decomposition (i.e., decomposing into row and column vectors) on the second-moment estimates of the gradients, Adafactor can reduce memory consumption from $O(m\cdot n)$ to $O(m+n)$. Meanwhile, Anil et al. proposed SM3 [114], which reduces the memory requirements for storing optimizer states through a parameter cover mechanism. It involves dividing parameters into multiple subsets, with each subset maintaining only one variable to approximate the second-order statistics of all parameters within it. This strategy helps decrease memory consumption during second-order moment estimation. SM3 employs a data-driven approach to dynamically adjust the learning rate, similar to Adagrad, while mitigating the memory overhead that arises from maintaining parameter-level statistics, as seen in Adagrad. Experiments demonstrate that the training performance of SM3 is on par with that of Adagrad and Adam, while its memory overhead is slightly lower than that of Adafactor.
Adafactor uses non-negative matrix factorization to approximate the second-order statistics of gradients. While it saves memory, it can introduce errors that lead to instability during training. SM3 encounters similar challenges. To tackle this issue, Luo et al. introduced CAME optimizer [115]. It utilizes the residual between the exponential moving average (EMA) of the update and the current update to compute a confidence matrix, subsequently adjusting parameter updates. By applying non-negative matrix factorization to the confidence matrix, CAME further decreases memory overhead. Through confidence-guided strategies and non-negative matrix factorization, CAME significantly reduces memory usage while ensuring convergence. In large-scale language model training tasks, CAME has shown superior performance compared to Adafactor, while being comparable to it in terms of memory efficiency.
Zhang et al. discovered that the Hessian matrix of the Transformer exhibits an approximate block diagonal structure. Based on this finding, they proposed Adam-mini [117], which partitions the parameters of components such as Q, K, V, and MLP into blocks and assigns a learning rate to each block, reducing the resource requirement by 90% to 99%. In the pre-training tasks of TinyLlama-1B, Llama2-7B, and the GPT series, Adam-mini demonstrated a significant reduction in memory overhead compared to AdamW, achieving savings of 45% to 50%. Additionally, it offered faster convergence speed while maintaining comparable or improved performance, which surpassed that of Adafactor, SM3, and CAME. Adam-mini effectively lowers the memory demands of the optimizer without compromising performance, providing a more efficient solution for training large-scale models.
IV Tailored Memory-Efficient Training Techniques: A Case Study

### v2

@@ 3.1.1 Compression
    Mixed Precision Training leverages data types of varying precisions to reduce memory consumption and computational costs, while preserving comparable accuracy. It was introduced by Micikevicius et al. [76] that model parameters, activations, and gradients are stored using FP16 while maintaining a copy of the FP32 parameters. To cope with numerical stability issues caused by reduced numerical range and accuracy, FP16 parameters are used for forward and backward calculations, while FP32 parameters are used for model updates. Additionally, to mitigate the risk of gradient underflow during backpropagation, loss scaling is commonly applied, which helps to minimize model performance degradation. The mixed precision training process is illustrated in Fig. 3.
@@
    While it needs to save an additional FP32 model parameter, the memory savings achieved by utilizing FP16 activations surpass the memory costs associated with these extra FP32 parameters, leading to an overall memory usage reduction.
@@
    Modern GPUs, such as NVIDIA’s Volta, Turing, and Ampere architectures, feature specialized Tensor Cores designed to accelerate mixed-precision matrix operations. However, BF16 support is limited to certain hardware platforms, such as Ampere and Hopper architectures. Since the Hopper architecture, FP8 Tensor Cores were added allowing for FP8 mixed precision training, which can further reduce memory consumption and accelerate computation [78, 79].
@@
    Mixed precision training requires careful management to prevent numerical issues. As hardware and software continue to evolve, the methods and tools for mixed precision training are also advancing, facilitating the training of larger and more intricate models.
  2. ii.
    Quantization-Aware Training
    Quantization-aware training (QAT) reduces memory consumption and computational demands by representing model parameters and activations in lower-precision formats (e.g., INT8 or FP8) while maintaining accuracy through training-time adaptation. Unlike post-training quantization (PTQ), QAT simulates quantization effects during forward and backward passes, enabling the model to learn robust representations resilient to precision loss.
    Liu et al. introduced LLM-QAT [80], a QAT framework for LLMs that requires no access to the original pre-training data. They generated synthetic data by sampling next-token outputs from the full-precision pre-trained model and trained quantized model via logit-based knowledge distillation. Applied to LLaMA-7B, 13B, and 30B, LLM-QAT achieves accurate 4-bit weight quantization and consistently outperforms SOTA PTQ methods, especially at precision lower than 8-bit. LLM-QAT can achieve 7.3× speedup in inference with 4-bit weight 8-bit activation quantization with comparable accuracy.
    BitNet [81] replaced standard linear layers with BitLinear modules using 1-bit weights and quantized activations. BitNet was trained via QAT and achieved competitive accuracy with significantly reduced memory and energy consumption compared to FP16 Transformers. The authors demonstrated that BitNet obeys a scaling law akin to full-precision Transformers, enabling efficient scaling to larger models while maintaining training stability and inference performance. Building upon this foundation, Ma et al. proposed BitNet b1.58 [82], a ternary-weight version (with weights in -1, 0, +1) that stroke a more favorable trade-off between accuracy and efficiency. They showed that b1.58-bit models matched or exceeded the performance of FP16 LLaMA models starting from 3B parameters while offering up to 4.1× decoding latency speedup, 7.2× memory savings, and 71.4× energy efficiency. Together, these works highlight the promise of low-bit quantization for scaling and deploying LLMs in a highly resource-efficient manner.
    EfficientQAT [83] aims to enhance memory and computational efficiency for LLMs while minimizing accuracy loss. It operates in two phases: Block-wise Training of All Parameters (Block-AP) and End-to-End Training of Quantization Parameters (E2E-QP). Experiment results showed that 2-bit EfficientQAT on Llama-2-70B completed training in 41h on a single A100-80GB GPU with peak memory $\approx$ 34 GB, achieving less than 3 points accuracy degradation compared to the FP16 model.
Table 2: An overview of algorithm-level memory-efficient training methods
  Category Method Feature Applicable Scenarios
Compression
Mixed Precision
Training
- Roughly halves overall memory consumption
- Speedup training
- May cause numerical issues
- Widely applicable
Quantization-Aware
Training
- Further compresses model size during training
- Prepares models for low-bit inference
- Potential loss in model accuracy
- Edge deployment scenarios
- Resource-constrained training
- When quantization compatibility is required
Memory Efficient
Optimizers
First-Order
- Fast convergence
- High memory demand
- Pre-training and finetuning
Zeroth-Order
- Low memory footprint
- Slow convergence
- Finetune on downstream tasks
- Resource constrained settings
Gradient Checkpointing
Full Checkpointing - Trades additional computation for reduced memory
- Avoid if memory is sufficient, as it decreases
throughput
Selective Checkpointing
- Perform recompute on components that cost substantial
memory yet require less computation
Gradient Accumulation -
- Increase effective batch size
- Delay weight updates across multiple steps
- Useful in low-memory environments
3.1.2 Memory Efficient Optimizers
Training LLMs often involve maintaining not only model parameters but also optimizer states such as momentum and variance estimates. These auxiliary states can triple the memory footprint, especially when using adaptive optimizers like Adam [117] or AdamW [118], which are widely adopted due to their fast convergence and training stability. However, this comes at a steep memory cost. Memory-efficient optimizers aim to reduce this overhead by redesigning the way these statistics are stored or approximated. These optimizers can be particularly beneficial in scientific applications where training resources are limited but model complexity is high.
Shazeer et al. introduced Adafactor [84], which significantly lowers memory requirements while delivering training performance comparable to Adam. Assuming the model parameter matrix is denoted as $W\in R^{m\times n}$, by applying low-rank decomposition on the second-moment estimates of the gradients, Adafactor can reduce memory consumption from $O(m\cdot n)$ to $O(m+n)$. Meanwhile, Anil et al. proposed SM3 [85], which reduces the memory requirements for storing optimizer states through a parameter cover mechanism. It involves dividing parameters into multiple subsets, with each subset maintaining only one variable to approximate the second-order statistics of all parameters within it. SM3 employs a data-driven approach to dynamically adjust the learning rate, similar to Adagrad, while mitigating the memory overhead that arises from maintaining parameter-level statistics, as seen in Adagrad. Experiments demonstrate that the training performance of SM3 is on par with that of Adagrad and Adam, while its memory overhead is slightly lower than that of Adafactor.
Adafactor uses non-negative matrix factorization to approximate the second-order statistics of gradients. While it saves memory, it can introduce errors that lead to instability during training. SM3 encounters similar challenges. To tackle this issue, Luo et al. introduced CAME optimizer [86]. It utilizes the residual between the exponential moving average (EMA) of the update and the current update to compute a confidence matrix, subsequently adjusting parameter updates. By applying non-negative matrix factorization to the confidence matrix, CAME further decreases memory overhead. Through confidence-guided strategies and non-negative matrix factorization, CAME significantly reduces memory usage while ensuring convergence. In LLM training tasks, CAME has shown superior performance compared to Adafactor, while being comparable to it in terms of memory efficiency.
Zhang et al. discovered that the Hessian matrix of the Transformer exhibits an approximate block diagonal structure. Based on this finding, they proposed Adam-mini [88], which partitions the parameters of components such as Q, K, V, and MLP into blocks and assigns a learning rate to each block, reducing the resource requirement by 90% to 99%. In the pre-training tasks of TinyLlama-1B, Llama2-7B, and the GPT series, Adam-mini demonstrated a significant reduction in memory overhead compared to AdamW, achieving savings of 45% to 50%. Additionally, it offered faster convergence speed while maintaining comparable or improved performance, which surpassed that of Adafactor, SM3, and CAME.
Traditional gradient-based optimizers update model parameters through gradients calculated during backpropagation, inevitably leading to additional memory overhead and computational complexity. On the other hand, zeroth-order optimizers estimate gradients based on function values, which can effectively reduce memory usage during model training. MeZO [89] was the first to use a zeroth-order optimizer for fine-tuning memory optimization of LLMs. By introducing a memory-efficient zeroth-order optimizer, the fine-tuning of LLM can be achieved with only forward passes, significantly lowering memory demands. As a comparison, MeZO can train a 3B model on a single A100 80GB GPU, while traditional optimizers can only train a 270M model under the same hardware budget.
DeepZero [90] primarily addresses the scalability issue of zeroth-order optimization in the training of DNNs. It highlights for the first time the benefits of coordinate gradient estimation compared to random vector gradient estimation. To enhance the deployment and application of zeroth-order optimizer training, DeepZero has introduced methods such as feature reuse and forward parallelization. On the other hand, Zo-AdaMU [91] handles the problems of oscillation, overfitting, and slow convergence in MeZO by incorporating momentum and uncertainty into simulated perturbations to refine gradient estimation. It also dynamically adjusts uncertainty and momentum parameters using a simulated annealing mechanism, thereby improving the stability and convergence speed of the optimization process.
3.1.3 Gradient Checkpointing
@@
During the training process of DNNs, the output of each layer (activations) needs to be stored in memory to calculate gradients during backpropagation. When dealing with large-scale transformers, long input sequences, or large batch sizes, the memory overhead caused by activations can become substantial. Gradient checkpointing reduces memory usage by retaining only the intermediate activations of certain layers during forward propagation while recalculating the discarded activations when needed during backpropagation. The core idea behind it is to trade computation for memory. Fig. 4 shows the difference in whether gradient checkpointing is enabled during the training process. The preserved intermediate activations are referred to as “checkpoints” [92].
@@
The number of parameters and computational complexity of various components in DNNs often vary. Korthikanti et al. [93] suggest that for transformer models, it is unnecessary to set checkpoints for the entire transformer layer for recomputation. Instead, selective recomputation can be performed by identifying components of the transformer layer where activations cost substantial memory yet require less computation (e.g., the computation of $Q\times K^{T}$, subsequent softmax, dropout, and attention score multiplies $V$). By setting checkpoints and recalculating selectively, memory consumption can be optimized without incurring significant computational overhead.
@@
3.1.4 Gradient Accumulation
When training large models, using a larger batch size is often desirable as it can improve training stability, convergence efficiency, and hardware utilization. However, GPU memory limitations may become a bottleneck. Gradient accumulation is helpful when GPU memory is insufficient to process large batches directly. The fundamental idea behind gradient accumulation is to gradually accumulate gradients from multiple small batches until a desired batch size is achieved, at which point a weight update is performed. For instance, if the total batch size we aim to use is 32, which exceeds the GPU memory capacity, we can divide it into smaller batches with batch sizes of 8, and perform 4 training iterations before conducting a weight update. Algorithm 1 presents the pseudocode of gradient accumulation, outlining the implementation of this technique in the training process.
3.1.5 Approximate Attention
While most memory-efficient techniques discussed in this survey focus on training-level optimizations, another important class of approaches aims to reduce memory complexity at the architectural level.
Linformer [119] reduces attention complexity from $O(n^{2})$ to $O(n)$ by projecting the key and value matrices into a lower-dimensional space using learned linear projections. This low-rank approximation significantly reduces both memory and compute cost with on-par model quality.
Performer [120] approximates softmax attention through kernel methods, replacing the softmax operation with a feature map transformation. It introduces the FAVOR+ (Fast Attention Via positive Orthogonal Random features) mechanism, enabling linear-time and linear-memory attention. The method maintains high accuracy while improving scalability to longer contexts.
These architectural modifications offer an orthogonal path to memory efficiency compared to the training-time techniques. Although such approximate attention models have not yet been widely applied in AI for Science tasks, they hold strong potential for future adoption. Since the primary focus of this survey is on training-time memory optimization, we only provide a brief overview of these methods here.
3.2 System
System-level techniques address memory limitations by leveraging inter-device communication and heterogeneous memory hierarchies. This section reviews representative system-level methods that support the training of transformer models through coordinated memory management across computing infrastructure.
3.2.1 Distributed Training
(a) Parameter Server
(b) Decentralized
Figure 5: Different architectures for data parallelism.
As the scale of models continues to increase, a single GPU’s memory can no longer accommodate complete model parameters, optimizer states, and intermediate activations. Traditional training strategies for large-scale models may take weeks or even months. However, distributed training allows the model to be divided across multiple devices, enabling the training of extremely large models with limited single GPU memory and significantly reducing training time. Table 3 provides a brief overview of the features, applicable scenarios, and performance of different distributed training techniques discussed in this section.
Table 3: An overview of distributed training methods
  Method Feature Applicable Scenarios Performance
DP DDP
- Accelerate training with larger global batch size
- moderate communication, can overlap with computation
- high redundancy
- Train models that fits in single GPU
(e.g., <1B params)
- Scale training to more GPUs / nodes
Training throughput scales linearly as the number of GPUs
increases
ZeRO-1
- partitioned optimizer states
- no extra communication
- redundant parameters and gradients
- Choose appropriate method according
to hardware environment
- Train larger models with comparable
throughput
- Combine with TP/PP to further expand
model size
4× memory redunction
ZeRO-2
- partitioned optimizer states and gradients
- no extra communication
- redundant parameters
8× memory reduction
ZeRO-3
- no redundancy
- 1.5× communication
$N_{d}$× memory reduction
TP
Megatron-LM
(1D-TP)
- Split matrix multiplication to reduce computation and
parameters stored on each device
- Frequent communication
- Redundant input activations
- High-speed bandwidth between adjacent GPUs
(e.g., NVLink)
- Apply within nodes to prevent communication
bottleneck
$N_{d}$× memory reduction, 77% throughput (8 GPUs)
2D-TP
- Lower communication volume
- More process groups
- Split input data
- GPUs interconnected with different
bandwidth
- Number of GPU >= 4
70% memory consumption when training with 4 GPUs
(compared to 1D-TP)
2.5D-TP
56% memory consumption when training with 8 GPUs
(compared to 1D-TP)
3D-TP
35% memory consumption when training with 8 GPUs
(compared to 1D-TP)
PP GPipe
- Partition model layers
- High peak memory usage
- Apply PP across nodes to scale up to larger
models
- Select methods based on actual needs
$N_{d}$× memory reduction, 3.5× throughput
(compared to vanilla PP)
1F1B - Lower peak memory usage
Reduced activation memory usage with comparable throughput
(compared to GPipe)
Interleaved 1F1B
- Finer-grained scheduling to further reduce bubbles
- More communication and activation memory usage
Throughput increased by 15% with 1.23× memory
(compared to 1F1B)
Zero Bubble
- Split backward pass
- Minimum bubbles
Throughput increased by 11% with 1.08× memory
(compared to 1F1B)
SP RSA - Ring-style calculation of self-attention - Incompatible with memory-efficient attention
71% memory consumption, 97% throughput
(compared to 1D-TP, sequence length = 2048)
Megatron SP - Divide layernorm and droupout - Apply in conjunction with Megatron-TP
67% memory consumption, 6% speedup
(compared to 1D-TP)
Ulysses - Utilize All-to-All communication to obtain attention heads
- Number of attention heads is an integer
multiple of SP degree
1.51× throughput vs. Megatron SP, 1.99× throughput vs. RSA
(sequence length = 8k, 32 GPUs)
Ring Attention
- Ring-style attention computation
- Communication can overlap with computation
- Applicable for both training and inference
Enable training large models (7B-65B) on long input (over 4M)
with comparable model FLOPS utilization
  1. i.
    Data Parallelism
[…]
    Inspired by ZeRO, PyTorch developed Fully Sharded Data Parallel (FSDP) [47]. It demonstrates performance comparable to that of DeepSpeed ZeRO, offering a robust solution for scaling deep learning models. Since PyTorch v2.4 , the PyTorch team has restructured FSDP and introduced FSDP2, adding additional optimizations to enhance performance.
  2. ii.
    Tensor Parallelism
    The rapid increase in model parameters has rendered the memory and computing resources of a single device insufficient to support the training of the entire model. Although ZeRO can partition the model across computing devices, it does not reduce computational workloads on each device. For large models with vast parameters and high computational demands, relying solely on DP becomes inadequate for meeting the training needs. Model parallelism (MP) has emerged for further scaling of models. MP primarily includes two techniques: tensor parallelism (TP) and pipeline parallelism (PP), which will be discussed in this section and the following section. TP, also referred to as intra-layer parallelism, finely divides the computation within each model layer into smaller tensor operations that can be executed concurrently across multiple devices, thereby effectively distributing both memory overhead and computational workload.
    Figure 6: Splitting of matrix multiplication.
[…]
    Megatron-LM TP divides tensors into one-dimensional segments. While this approach reduces model parameters stored on each device to $\frac{1}{N_{d}}$, it still needs to store complete intermediate activations on each device. Additionally, forward and backward computations of each transformer layer involve 4 all-reduce, causing significant communication overhead. To address this, researchers have proposed multidimensional tensor partitioning methods [97, 98, 99], which further reduce memory overhead on each device and mitigate the frequent communication issues of 1D TP [123]. In practice, TP is typically applied within nodes rather than across nodes to prevent communication from becoming a bottleneck for model training efficiency [93].
  3. iii.
    Pipeline Parallelism
    Pipeline parallelism, or inter-layer parallelism, involves partitioning the model layers horizontally into multiple stages. Each stage is assigned to different devices for sequential execution. PP not only reduces memory demands and computational workloads on each device but also has less communication overhead. However, a major challenge associated with PP is device idle time (often referred to as pipeline bubbles). While asynchronous PP schemes [101, 124] can eliminate bubbles, they have no guarantee of model convergence [105]. As synchronous PP schemes continue to improve, their efficiency are comparable with asynchronous PP, making them a more favorable option.
    (a) Pipeline schedules (top: 1F1B, mid: interleaved 1F1B, bottom: ZB-H1, $D_{i}$ represents device $i$)
    (b) Device placements (left: 1F1B / ZB-H1, right: interleaved 1F1B)
    Figure 7: Pipeline schedules and device placements of 1F1B, interleaved 1F1B, and ZB-H1. Interleaved 1F1B partitions model into finer-grained stages and assign each device with multiple stages to reduce pipeline bubbles.
    In vanilla PP, data can be passed to the next device only after the preceding device completes its calculation. This leads to a waste of resources and diminishes system efficiency. GPipe [100] splits each input batch into several micro-batches. After completing the processing of each micro-batch, each device immediately passes its results to the next device. GPipe significantly reduces pipeline bubbles, however, it faces the issue of high activation memory consumption.
    Let $N_{s}$ denote the number of stages in the model, $N_{d}$ denote the number of devices, and $N_{m}$ denote the number of micro-batches. In GPipe, each device must retain activations of $N_{m}$ micro-batches. To reduce memory consumption, researchers introduced the 1F1B schedule [101, 102, 103] (see Fig. 7a). This schedule effectively reduces the number of on-the-fly micro-batches by executing the backward pass in advance. Its pipeline bubble is equal to that of GPipe, and the activations of micro-batches required to store on each device decrease from $N_{m}$ to $N_{d}$.
    The aforementioned PP schemes assume that $N_{s}=N_{d}$, with each device possessing one stage. Narayanan et al. [102] split the stages, allowing each device to have multiple stages (we demonstrate an example in Fig. 7b). This interleaved 1F1B schedule (Fig. 7a) performs finer-grained scheduling, reducing pipeline bubbles and enhancing efficiency. However, the trade-off for this includes increased communication and higher activation memory usage.
    The computational cost of backpropagation is greater than that of forward pass, resulting in longer computation time, causing the existence of pipeline bubbles. Qi et al. [104] suggest splitting the backward pass into calculating the gradients of activations and the gradients of parameters (denote the two calculations as $B$ and $W$ respectively). Theoretically, one could complete $B$ passes of all layers before performing $W$ passes for each layer. Based on this finding, they proposed ZB-H1 schedule, which offers the flexibility to schedule $W$ after the corresponding $B$ within the same stage, effectively filling bubbles in the schedule (depicted in Fig. 7a).
    PP typically has the issue of high activation memory usage, and the 1F1B schedules mentioned above have the problem of memory imbalance. Qi et al. [105] proposed an analytic framework to decompose existing pipeline schedules into building blocks and found that peak memory usage of a pipeline schedule is highly related to the lifespan of the building blocks.
    Compared to TP, PP only communicates between adjacent stages and has lower communication overhead. Therefore, it is usually applied between nodes to scale up model size [102].
  4. iv.
    Sequence Parallelism
    Supporting long sequence lengths is crucial for applying AI to scientific endeavors. Many scientific challenges inherently involve complex analysis of large-scale, high-dimensional data, frequently presented as long sequences. Long-sequence issues are prevalent in fields including structural biology, chemistry, drug development, and atmospheric science. For instance, many proteins consist of hundreds or even thousands of amino acid residues, where the sequence context of each residue can be essential to its 3D structure. Long-sequence support enhances the ability to capture non-local interactions among residues, thereby improving the quality of predictions. Additionally, complex molecules or chemical reactions often necessitate long sequence representations, including molecular formulas, SMILES, reaction pathways, etc. Climate prediction models require analyzing long-term historical and global spatial data, represented in time series or grid formats. Despite the importance of long-sequence support, processing long sequences demands significant increases in computational power and memory, with the memory usage of attention mechanisms rising quadratically with sequence length.
    Figure 8: Workflow of Ulysses sequence parallelism with 2 devices. Notations: $s$ for sequence length, $h$ for hidden size, $a$ for number of attention heads, and $p$ for SP degree.
    Figure 9: Workflow of Ring Attention sequence parallelism with 2 devices.
    Jacobs et al. proposed Ulysses [107], a method that divides the input along sequence dimension onto all devices and conducts all-to-all communication on the Query (Q), Key (K), and Value (V) matrices before performing attention computation. This ensures that each device operates with disjoint attention heads, allowing for distributed attention calculation. Once the calculations are complete, another all-to-all is performed to get the results and restore the state that inputs were divided along sequence dimension. This process is presented in Fig. 8. Ulysses SP and Megatron-LM TP share similarities in the distributed calculation of self-attention, both concerning limitations on the number of attention heads. A key advantage of Ulysses is that the communication volume for a single device decreases linearly as SP degree increases, while in Megatron-LM, it is independent of TP degree.
    Liu et al. [108] developed a ring-style distributed attention computation method that optimizes resource usage by conserving a portion of the Q, K, and V blocks. During calculations, each device exchanges K and V blocks with its neighboring devices, enabling it to perform computations using blockwise self-attention and feedforward networks (FFN) [125], as depicted in Fig. 9. In contrast to the approach taken by Li et al. [106], this method can eliminate communication overhead by picking a proper chunk size and overlapping communication with computation.
3.2.2 Offloading
SwapAdvisor [109] takes the data flow graph of a model as input and selects legitimate memory allocation and operator scheduling based on the graph. By precisely planning which and when tensors require swapping in and out, SwapAdvisor achieves optimal overlap between computation and communication. The integration of genetic algorithms enables the exploration of the vast search space for finding the optimal combination of memory allocation and operator scheduling. In contrast to previous heuristic-based swapping methods [126, 127], SwapAdvisor significantly enhances memory efficiency and computational performance through its dual optimization approach. Notably, due to the lack of consideration of GPU communication, SwapAdvisor has limitations in multi-GPU training scenarios.
ZeRO-infinity [111] enhances the capabilities of ZeRO-3 by adding support for heterogeneous memory, thereby leveraging the entire spectrum of memory types within the system (GPU, CPU, and SSD) to enable large-scale model training. By employing the Infinity Offload Engine to offload model states to NVMe storage and efficient communication optimizations, ZeRO-infinity enables super-linear scalability of training throughput when scaling from 64 GPUs to 512 GPUs for training a 1T-parameter model, the overall throughput increased from 2.8 petaflops to 25 petaflops, surpassing linear scaling. ZeRO-Infinity trains 500B model with throughput comparable to 3D-parallelism, if increasing the model size to 1T, 3D-parallelism runs out of memory, while ZeRO-Infinity still reaches over 70% of achievable hardware FLOPS utilization. On 32 Nvidia V100 DGX-2 nodes, ZeRO-infinity supports Transformer models with up to 32T parameters, which is 50× the maximum size accommodated by 3D-parallelism (around 650B parameters).
A comparative summary of representative offloading strategies, including SwapAdvisor, ZeRO-Offload, and ZeRO-Infinity, is provided in Table 4, highlighting their core features, applicable scenarios, and performance.
Table 4: An overview of offloading methods
  Method Feature Applicable Scenarios Performance
SwapAdvisor
- Optimize scheduling, memory allocation
and swap planning simultaneously
- Single-GPU training
Achieve 53% - 99% of the training throughput of the
ideal baseline
ZeRO-Offload - Utilize CPU to perform weight update
- Apply in conjunction with ZeRO-2 and
mixed precision
Outperforms SwapAdvisor by 23% in training throughput;
Enables training 1.62× larger model (vs. SwapAdvisor) on a
single GPU, and increases model scale by 4.5× (vs. Megatron)
and 7.8× (vs. ZeRO-2) on a single DGX-2 node
ZeRO-Infinity - Exploit CPU and NVMe memory
- Apply in conjunction with ZeRO-3
- Train extremely large models
(e.g., >1T params)
Enable 1T model training on a single DGX-2 node without MP
3.3 Hardware-Software Co-Optimization
As transformer models scale, training them efficiently requires both algorithmic innovations and tight coupling between hardware characteristics and software design. By aligning software execution patterns with hardware capabilities—such as memory hierarchy, bandwidth, and parallelism—these techniques significantly improve training throughput and memory efficiency.
A representative example of hardware-software co-optimization in training transformers is FlashAttention [112], which re-architects the attention computation to exploit GPU memory hierarchy, achieving significant speed and memory efficiency improvements. It introduces an IO-aware exact attention algorithm that significantly improves memory efficiency by minimizing expensive reads/writes between high-bandwidth GPU memory (HBM) and on-chip SRAM. Through tiling and recomputation, FlashAttention avoids materializing large intermediate attention matrices, achieving up to 3.5× training time speedup and reducing memory consumption of self-attention from $O(n^{2})$ to $O(n)$, which enables scaling transformer models to longer sequences.
FlashAttention-2 [113] builds upon FlashAttention by optimizing GPU work partitioning and parallelism. It reduces non-matrix-multiply operations, increases GPU occupancy by parallelizing across sequence dimensions, and minimizes shared memory communication. As a result, FlashAttention-2 achieves 2× speedup over FlashAttention and reaches up to 73% of the theoretical peak FLOPS on A100 GPU. Focusing on hardware-aware optimizations for Hopper architecture, FlashAttention-3 [114] optimizes attention computation on GPUs using asynchronous execution, interleaved GEMM-softmax operations, and FP8 low-precision with block quantization. It achieves 1.5–2.0× speedup over FlashAttention-2 (85% utilization in BF16, 1.3 PFLOP/s in FP8) while reducing FP8 error by 2.6× compared to baseline FP8 attention.
MPress [115] is a novel system designed to break the GPU memory wall for billion-scale model training on multi-GPU servers. It combines inter-operator parallelism with a new D2D swap technique to efficiently transfer tensors between GPUs, leveraging high-bandwidth NVLink connections. MPress dynamically selects the best memory-saving strategies based on tensor live intervals and available spare memory. Evaluations show that MPress can train larger models compared to recomputation baseline (3.7× for BERT and 1.7× for GPT) while maintaining high training throughput, and achieves 1.4–2.3× speedups compared to ZeRO-Series baselines [110, 111] under the same memory reduction.
Beyond these optimizations, broader efforts in the LLM community have also explored general-purpose hardware–software co-design frameworks. Wattanawong et al. [128] propose an approach that jointly evolves Transformer architectures and hardware accelerators via neural architecture search. Their framework achieves up to 10× improvement in energy-delay product (EDP) with minimal perplexity degradation, demonstrating that hardware-aware architectural adaptations can substantially reduce inference costs. Meanwhile, Guo et al. [129] present a comprehensive survey of LLM-centric co-design, outlining the challenges posed by large model size, energy consumption, and memory bottlenecks.
4 Tailored Memory-Efficient Training Techniques: A Case Study